OpenStudy (anonymous):

Simplify by using radical. Rationalize denominators.. 2a^-1/2. Please please help!!

OpenStudy (campbell_st):

well you have \[\frac{2}{\sqrt{a}}\] just multiply by 1.... you just need to write it in an appropriate form... such as \[\frac{1}{1} = \frac{\sqrt{a}}{\sqrt{a}}\]

OpenStudy (anonymous):

Thank you so much! can you help me with a few more?

OpenStudy (campbell_st):

just post then in the normal way... there are lots of people who can help